# Cosmology from point clouds

### v1

@@ Abstract
We train a novel deep learning architecture to perform likelihood-free inference on the value of the cosmological parameters from halo catalogs of the Quijote N-body simulations. Our model takes as input a halo catalog where each halo is characterized by its position, mass, and velocity moduli. By construction, our model is E(3) invariant and is designed to extract information hierarchically. Unlike graph neural networks, it does not require the transformation of the input halo (or galaxy) catalog into a graph. Given its simplicity, our model can process point clouds with large numbers of points. We discuss the advantages of this class of methods but also point out its limitations and potential ways to improve them for cosmological data.
@@ 1 Introduction
Our ability to characterize and understand our Universe depends on the level of accuracy to which the value of the different cosmological parameters can be constrained. Since the spatial distribution of galaxies in the sky is sensitive to the value of different cosmological parameters, it has been a common practice to use galaxy clustering to learn about the laws and constituents of the Universe (Uhlemann et al., 2020; Villaescusa-Navarro et al., 2020; Gualdi et al., 2021; Valogiannis & Dvorkin, 2021; Liu et al., 2022; Ajani et al., 2020), especially with upcoming sky surveys like CMB-S4111https://cmb-s4.org/, Roman [^2] https://roman.gsfc.nasa.gov/, Dark Energy Spectroscopic Instrument (DESI) (DESI Collaboration et al., 2016), Euclid (Laureijs et al., 2011; Racca et al., 2016; Euclid Collaboration: Castro et al., 2022), the Large Synoptic Survey Telescope (LSST) (Ivezić et al., 2019), J-PAS (Benitez et al., 2014), that are expected to collect data with unprecedented accuracy, this method will be extremely useful for precision cosmology.
@@
In this work, we employ a novel architecture that takes a point cloud as input and processes it similarly to graph neural networks (GNNs). The advantage of this model is that it does not require an input graph (as GNNs do) and can also extract information on different scales hierarchically, allowing the processing of relatively large point clouds. We note that some previous works exploited complicated geometrical feature extractors to process point clouds (Li et al., 2019; Guo et al., 2020). Recently, Ma et al. (2022) proposed a novel method that takes point clouds as input and processes them in a very similar fashion as GNNs (i.e., with simple, flexible, and learnable geometrical kernels) without having to predefine a graph from the point cloud. They have shown that their method either outperforms or performs at par with other state-of-the-art point cloud networks. Further, due to employing a simple feature extractor, this network is lightweight and fast compared to other 3D data analysis methods.
@@ 2 Data
In this paper, we work with 3D point clouds. A point cloud contains N points, and each point is characterized by its 3D position (x,y,z). The points can also have other properties, such as masses and peculiar velocity moduli. Each point cloud is fully described by five labels: the values of the cosmological parameters. In this work, we will, however, restrict ourselves to just two of these parameters: $\Omega_{\rm m}$ and $\sigma_{8}$.
@@ 3.1 Architecture
For each simulation, we have a collection of $N$ points $\{p_{1},p_{2},...p_{n}\}$ where each point represents a dark matter halo. We will denote the properties of the point $i$ by $f_{i}\in\mathbb{R}^{d}$. The properties of the $j$-th closest neighbour of the point $i$ are instead represented by $f_{i,j}\in\mathbb{R}^{d}$. The dimensionality of the input feature vector is
@@
  • Stage. At the beginning of a stage, a subset of $n$ points is selected using the farthest point sampling algorithm. In our case, we select half of the input points. Next, the features of the selected point are updated as follows.
@@
    $\delta=10^{-5}$ is added for numerical stability. Next, we update the features of point $i$ using the transformed properties of its $k$ nearest neighbors via
@@ 3.2 Symmetries
The values of the cosmological parameters of a set of dark matter halos are not affected if we translate and/or rotate their positions. These symmetries can be learned in our model through data augmentation or imposed directly into the model. In our case, we decide to impose them using invariant input features of the point cloud.
@@ 3.3 Loss Function
where $\theta_{i,j}$, $\mu_{i,j}$, $\sigma_{i,j}$ represent the true, inferred mean, and inferred standard deviation of the parameter $i$ for the sample $j$.
@@ 4 Results
Table 1: This table summarizes the main results of this work. The first column states the configuration used with the notation $N_{k}$, where $N$ is the number of halos in the catalog and $k$ is the number of neighbors. The second column shows the features used for the halos, where pos, M, and V stand for positions, masses, and velocity moduli. 2ptCF denotes the usage of the 2-point correlation function using just the positions of the dark matter halos. The other columns show the value of the different validation metrics for $\Omega_{\rm m}$, and $\sigma_{8}$.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 5 Model limitations and improvements
  • In our model, we use max-pooling layers to update the features of the selected points from their neighborhood. It will be interesting to study whether other aggregation schemes (e.g., sum or mean) or a combination of them (Villanueva-Domingo, 2022) yield better results.
  • For point clouds with a large number of points, we may need to make our model deeper, i.e., to include more stages. This will allow clustering information to be captured in a hierarchical manner and perhaps facilitate the work of the inference layer.
  • We have used the farthest point sampling algorithm to subsample the number of points in each stage. It is unclear whether that method is the best one to use in cosmology. It will be interesting to try alternative methods (e.g., random subsampling) that may better preserve the information from large overdense regions and penalize underdense regions.
@@ 6 Conclusions
The model explored in this work presents several advantages with respect to previous methods (e.g., message-passing GNNs). First, its simplicity and lack of sophisticated geometric kernels make the model fast to evaluate. Second, the model can work with large point clouds. In this work, we have explored models that contain up to 8,192 points, but this number can easily made much larger if needed (e.g., decreasing the number of points to sample during the stages) [^8] Without the requirements for periodic boundary conditions, we have been able to train models containing up to 80,000 points with a relatively large batch size. Thus, for lightcone datasets, our model can easily deal with hundreds of thousands of points if several GPUs are used.. Third, its simplicity makes it very easy to adapt to different tasks. Fourth, this model works with the natural representation of the data (a point cloud), while GNNs need a graph (it is unclear how to construct the optimal graph from a point cloud) to operate. We note that deep sets are deep learning architectures designed to work with data that can be represented as sets, but in general, they do not exploit the spatial information (see e.g. Wang et al., 2023, for an example of using deep sets to infer cosmology from void catalogs).
@@
Finally, we note that while in this work we have worked with halo catalogs from N-body simulations, the method can be applied to other point cloud data, like galaxy catalogs from hydrodynamic simulations or semi-analytic models (Villaescusa-Navarro et al., 2022b; Perez et al., 2022) or void catalogs (Kreisch et al., 2022).
Acknowledgements
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption

### v2

@@ Abstract
We train a novel deep learning architecture to perform likelihood-free inference on the value of the cosmological parameters from halo catalogs of the Quijote N-body simulations. Our model takes as input a halo catalog where each halo is characterized by its position, mass, and velocity modulus. By construction, our model is E(3) invariant and is designed to extract information hierarchically. Unlike graph neural networks, it does not require the transformation of the input halo (or galaxy) catalog into a graph. Given its simplicity, our model can process point clouds with large numbers of points. We discuss the advantages of this class of methods but also point out its limitations and potential ways to improve them for cosmological data.
@@ 1 Introduction
Our ability to characterize and understand our Universe depends on the level of accuracy to which the value of the different cosmological parameters can be constrained. The spatial distribution of galaxies in the sky is influenced by various cosmological parameters. As a result, galaxy clustering has become a common practice to study the laws and components of the Universe (Uhlemann et al., 2020; Villaescusa-Navarro et al., 2020; Gualdi et al., 2021; Valogiannis & Dvorkin, 2021; Liu et al., 2022; Ajani et al., 2020). Upcoming sky surveys—such as CMB-S4 (Abazajian et al., 2022), Roman Space Telescope (Lam et al., 2023), Dark Energy Spectroscopic Instrument (DESI) (DESI Collaboration et al., 2016), Euclid (Laureijs et al., 2011; Racca et al., 2016; Euclid Collaboration: Castro et al., 2022), the Large Synoptic Survey Telescope (LSST) (Ivezić et al., 2019), J–PAS (Benitez et al., 2014) will gather data with unprecedented precision. This will make galaxy clustering an especially powerful method for advancing precision cosmology.
@@
In this work, we employ a novel architecture that takes a point cloud as input and processes it efficiently using Multi-Layer Perceptrons (MLPs), whereas Graph Neural Networks (GNNs) rely on graph-based structures and message passing. A key advantage of this model is its ability to extract information hierarchically, enabling the efficient processing of relatively large point clouds. We note that some previous works exploited complicated geometrical feature extractors to process point clouds (Li et al., 2019; Guo et al., 2020). Recently, Ma et al. (2022) proposed a novel method that takes point clouds as input and processes them in a very similar fashion as GNNs (i.e., with simple, flexible, and learnable geometrical kernels) without having to predefine a graph from the point cloud. They have shown that their method either outperforms or performs at par with other state-of-the-art point cloud networks. Further, due to employing a simple feature extractor, this network is lightweight and fast compared to other 3D data analysis methods.
@@ 2 Data
In this paper, we work with 3D point clouds. A point cloud contains N points, and each point is characterized by its 3D position (x,y,z). The points can also have other properties, such as masses and peculiar velocity moduli  [^2] This is defined as $v=\sqrt{v_{x}^{2}+v^{2}_{y}+v^{2}_{z}}$ where $v_{x}$, $v_{y}$, $v_{z}$ are the cartesian components of the peculiar velocity respectively. Each point cloud is fully described by five labels: the values of the cosmological parameters. In this work, we will, however, restrict ourselves to just two of these parameters: $\Omega_{\rm m}$ and $\sigma_{8}$.
@@ 3.1 Architecture
For each simulation, we have a collection of $N$ points $\{p_{1},p_{2},...p_{n}\}$ where each point represents a dark matter halo. We will denote the properties of the point $i$ by $f_{i}\in\mathbb{R}^{d}$. The properties of the $j$-th closest neighbor of the point $i$ are instead represented by $f_{i,j}\in\mathbb{R}^{d}$. The dimensionality of the input feature vector is
@@
  • Stage. At the beginning of a stage, a subset of $n$ points is selected using the Farthest Point Sampling (FPS) algorithm. In our case, we select half of the input points. Next, the features of the selected point are updated as follows.
@@
    $\delta=10^{-5}$ is added for numerical stability. As mentioned in Ma et al. (2022) the affine transformation is introduced for transforming the local points to a normal distribution which in turn makes the network more robust, stable and accurate. Next, we update the features of point $i$ using the transformed properties of its $k$ nearest neighbors via
@@ 3.2 Symmetries
The values of the cosmological parameters of a set of dark matter halos are not affected if we translate and/or rotate their positions. These symmetries can be learned in our model through data augmentation or imposed directly into the model. In our case, we decide to impose them using invariant input features of the point cloud. We checked that the data augmentation (random rotation, translation etc) increases the training time of the model significantly, where as making the input features E(3) invariant bypasses the need for any data augmentation and makes the training much faster.
@@ 3.3 Loss Function
where $\theta_{i,j}$, $\mu_{i,j}$, $\sigma_{i,j}$ represent the true, inferred mean, and inferred standard deviation of the parameter $i$ for the sample $j$. The reason for using this loss function is that it bypasses the need to compute the posterior density of each parameter, directly obtaining the estimates of moments of the marginalized distribution of all the parameters.
@@ 4 Results
$8192$ Pos+M+V (XgBoost) 11.4 0.83 - $1.8\times 10^{-3}$ 5.2 0.76 - $2.6\times 10^{-3}$
Table 1: This table summarizes the main results of this work. The first column states the configuration used with the notation $N_{k}$, where $N$ is the number of halos in the catalog and $k$ is the number of neighbors. The second column shows the features used for the halos, where pos, M, and V stand for positions, masses, and velocity moduli. While the 2ptCF denotes the usage of the 2-point correlation function using just the positions of the dark matter halos, XGBoost indicates the cases where the XGBoost model are used. The other columns show the value of the different validation metrics for $\Omega_{\rm m}$, and $\sigma_{8}$.
@@ 4.1 Comparison with 2 point correlation function
4.2 Comparison with XGBoost
We also benchmark our study with XGBoost [^7] https://xgboost.readthedocs.io/en/stable/index.html (Chen & Guestrin, 2016), a simple non-neural random-forest based algorithm. The aim of this exercise is to highlight how a random-forest based regression algorithm with Pos+M+V features performs compared to PointMLP-elite. A key distinction between XGBoost and PointMLP-elite is that the former relies solely on the Pos+M+V features of the selected halos, without incorporating any information about the underlying topology from the k-nearest neighbors. We also note that XGBoost can only estimate the mean values of the cosmological parameters but not the standard deviation, therefore, we could not provide the $\chi^{2}$ values in this case.
We have trained the XGBoost model with $\rm n_{estimators}=400$ (size of the forest to be trained), max_depth $=10$ (maximum depth of a tree), and tree_method= ‘hist’ (tree construction algorithm used in XGBoost) and presented the results in Table 1.
We find that the XGBoost model performs significantly worse compared to the PointMLP-elite network. Especially with 8192 halos, the mean squared errors obtained from XGBoost are nearly an order of magnitude higher than those from the PointMLP-elite. Additionally, the mean relative errors from XGBoost are twice as large as those achieved by PointMLP-elite. Therefore, it is evident that PointMLP-elite model has gained significant insight by learning the topology around a halo through feature extraction from its nearest neighbors.
4.2.1 Component Ablation Study
In this subsection, we have described in detail the ablation studies demonstrating how the model performs when we changed individual components in PointMLP-elite architecture. We have summarized the result of the ablation study in table-2 and described them as follows.
We have mentioned on the last row of Table-1, the values of different validation matrices for both $\Omega_{m}$, $\sigma_{8}$. To reiterate, we have obtained that result with (a) configuration $8192_{32}$, (b) features Pos+M+V, (c) FPS subsampling, (d) Max-pool aggregation, (e) 4 layers. For the purpose of the ablation study, we kept that configuration as our benchmark and then varied different individual parts of the PointMLP-elite network as follows
  1. 1.
    Network depth: We varied the number of stages from 3 to 6. As shown in Table-2, the accuracy of the model declined with 3 stages. The accuracy significantly increased with 4 stages, but then we found that the performance of the model did not improve after 4 stages; only the time required for execution of the codes increased.
  2. 2.
    Max Pooling: we have then tried changing the aggregation method from Max-Pooling to the mean and sum aggregation but found no significant change in the model performance.
  3. 3.
    Subsampling Method: We have also experimented with the subsampling method. We tried a random subsampling technique rather than the FPS, but the accuracy of the model remains unchanged.
Fixed part Varying part Features $\Omega_{m}$ $\sigma_{8}$
$\epsilon(\%)$ $R^{2}$ MSE $\epsilon(\%)$ $R^{2}$ MSE
Max-pool, FPS Number of Layers 3 layers 8.3 0.89 $2\times 10^{-3}$ 6.2 0.75 $8.9\times 10^{-3}$
4 layers 7.1 0.97 $4.6\times 10^{-4}$ 5.2 0.81 $2.6\times 10^{-3}$
5 layers 6.9 0.95 $4.9\times 10^{-4}$ 5.9 0.85 $2.4\times 10^{-3}$
6 layers 7.5 0.98 $4.7\times 10^{-4}$ 5.5 0.81 $2.3\times 10^{-3}$
4 layers, FPS Aggregation Max-pool 7.1 0.97 $4.6\times 10^{-4}$ 5.2 0.81 $2.6\times 10^{-3}$
Mean 7.5 0.94 $4.8\times 10^{-4}$ 5.0 0.79 $2.5\times 10^{-3}$
Sum 7.7 0.95 $5.0\times 10^{-4}$ 5.4 0.83 $2.7\times 10^{-3}$
4 layers, Max-pool Subsampling FPS 7.1 0.97 $4.6\times 10^{-4}$ 5.2 0.81 $2.6\times 10^{-3}$
Random 7.2 0.95 $4.2\times 10^{-3}$ 5.3 0.83 $2.9\times 10^{-3}$
Table 2: This table summarizes the results obtained by varying different hyperparameters/parts of the network. The first column states the parts of the architecture that were kept fixed, while the second column shows the varying parts of the network. The third column mentions the details of the varying features. For all the scenarios, we use 8192 halos with 32 nearest neighbors, i.e.,$8192_{32}$, with Pos+M+V as the input feature.
@@ 6 Conclusions
The model explored in this work presents several advantages with respect to previous methods (e.g., message-passing GNNs). First, its simplicity and lack of sophisticated geometric kernels make the model fast to evaluate. Second, the model can work with large point clouds. In this work, we have explored models that contain up to 8,192 points, but this number can easily be made much larger if needed (e.g., decreasing the number of points to sample during the stages).  [^8] We would like to emphasize that we dealt with 80,000 halos (when the periodic boundary condition is ignored) in this work which is not even feasible for GNN. Thus, for lightcone datasets, our model can easily deal with hundreds of thousands of points if several GPUs are used. This low computation time for PointMLP is extremely important in the context of future sky surveys. In the upcoming sky surveys, the number of the galaxy will be very large, and dealing with a large number of galaxies using GNN is not possible even with state-of-the-art computers. Therefore, a model like PointMLP with a computational time less than that of a GNN will be very important for future galaxy surveys. Third, its simplicity makes it very easy to adapt to different tasks. Fourth, this model works with the natural representation of the data (a point cloud), while GNNs need a graph (it is unclear how to construct the optimal graph from a point cloud) to operate. We note that deep sets are deep learning architectures designed to work with data that can be represented as sets, but in general, they do not exploit the spatial information (see e.g. Wang et al., 2023, for an example of using deep sets to infer cosmology from void catalogs).
@@
Finally, we note that while in this work we have worked with halo catalogs from N-body simulations, the method can be applied to other point cloud data, like galaxy catalogs from hydrodynamic simulations or semi-analytic models (Villaescusa-Navarro et al., 2022b; Perez et al., 2022) or void catalogs (Kreisch et al., 2022). The main reason for choosing the halo catalogs rather than the galaxy catalogs is the unavailability of training data sets with volumes comparable to the upcoming sky surveys. The state-of-the-art Hydrodynamical data suit available to us is CAMELS, which covers a volume of $(25h^{-1}Mpc)^{3}$, much smaller than the volume of future sky surveys. As this work is a precursor to those sky-surveys data, we chose Quijote simulation with a box size of 1 $h^{-1}$ Gpc, comparable to the upcoming sky surveys.
Acknowledgments